# Malsight: Exploring Malicious Source Code and Benign Pseudocode for Iterative Binary Malware Summarization

### v1

@@ 5.1 Evaluation Dataset Construction
Utilizing the code summary from the MalS dataset as a foundation for our research, we curated a positive and negative sample pair for the tuning of our new evaluation model BLEURT-sum. The positive sample consisted of two code summary result sentences sharing the same meaning and was initialized as $\{S_{g},S_{r},1\}$, while the negative sample comprised two randomly different sentences and was initialized as $\{S_{g},S_{r},0\}$. $S_{g}$ and $S_{r}$ represent generated statement and reference statement, respectively.
@@
In order to build the dataset in $\{S_{g},S_{r},Score\}({Score}\in[0,1])$ format, one possible idea is to build an algorithm, $Score=GenSim(S_{g},S_{r},0or1)$, to automatically generate the label $Score$ required for model training. Therefore, the key task is to construct a reliable $GenSim()$ function. When $S_{g}==SUM(f_{n})$, contrusted $Score$ can be expressed using equation (11), considering that the SUM() function produces different outputs when faced with the same input, which is a non-deterministic function.
@@
Since when $\{S_{g},S_{r}\}$ pairs were built, no sentence structure dependencies are taken into account, which means the equation (11) of Score can contain a few words overlap-based features (§ 2.3), the natural consideration is to combine the characteristics of sentence structure and semantics. Our idea is to solve the proportion of semantic information and sentence structure information in sentence similarity evaluation.
Taking the original 0,1 tag as the semantic feature, we further extract the static feature to get a Score, multiply the two by the corresponding proportion respectively, and then add them to get $Score$. Assuming that the proportion of semantic information is $p$, the structural feature calculation function is $Struc()$, the following equations formalize this Score construction method.
$s_{f}=Struc(S_{g},S_{g})$ (12)
@@
5.2 Evaluation Model
Based on our insights into the evaluation model, it should be able to combine both structural and semantic features and give an evaluation score for the specific task of code summary.
@@ 5.2 Evaluation Model
The rich training signals ensure the universality of BLEURT and the comprehensiveness of the evaluation angle. We then further fine-tune the BLEURT model on the code summary sentence pair dataset. First, the generated text and the reference text are input together into the model and the vector is generated as shown in equation (14). This solution is called BLEURT-sum.
@@
In existing works, Word overlap measures are commonly used in text similarity evaluation but perform poorly for code summary evaluation. Code summaries often include keywords like “retur” and “initializ”, which do not significantly contribute to the overall meaning. These keywords can inflate overlap rates and lead to misleadingly high scores. Therefore, BLEURT-sum on the one hand still considers Word overlap measures, and on the other hand introduces more dimensions, which greatly improves the performance in code summary evaluation tasks.
6 Experiments
Our evaluation experiment was designed to answer the following four questions:
  1. 1.
    RQ1 (§ 6.2). How does Malsight performance compare across two training phases and mainstream code summarization models?
  2. 2.
    RQ2 (§ 6.3). How do module combinations and different stripping scenarios affect Malsight’s performance, particularly the annotator module?
  3. 3.
    RQ3 (§ 6.4). How does the new BLEURT-sum evaluation method compare to existing code summarization evaluation metrics?
  4. 4.
    RQ4 (§ 6.5). How does Malsight perform when applied to real-world malicious software, and how do assessments from human reverse engineers validate its usability?
6.1 Experimental Setup
Our experiment is running on the Ubuntu 20.04 system, equipped with one Intel Xeon Silver 4210 CPU 2.20 GHz, and two NVIDIA A40 GPUs with 125 GB RAM. Binary file processing tools include Radare2, IDA Pro v7.5, GCC v9.4.0, and GNU Make v4.2.1. Our programming language is Python v3.8.13, with transformers v4.16.2, torch v2.1.2+cu121.
6.2 RQ1. Performance Test
6.2.1 Code Summary Performance
During training, the model undergoes two phases. The first phase involves fine-tuning malware source code (MalS) to learn malicious behavioral features. The second phase focuses on generating high-quality code summaries (BenignC) by incorporating semantic information from pseudocode and annotations. In both phases, we split the dataset into training, cross-validation, and test sets with a ratio of 7:2:1. We then evaluated the model at each stage to ensure its effectiveness. The result is shown in Table II.
Since each phase uses a different dataset, their effectiveness should be verified independently rather than by direct score comparison. Considering there is no universally accepted threshold for model evaluation scores, we provide the performance of another state-of-the-art binary code summarization model (CP-BCS [55]) as a baseline. CP-BCS was tested on its own dataset, which consists of binaries compiled with GCC 7.3.0 for x86 architecture (32-bit) and then stripped. Our results in both phases demonstrate higher usability compared to CP-BCS.
TABLE II: Score During Two Train Phase
BLEURT-sum BLEU ROUGE-L METEOR
Phase1 74.17 18.06 38.56 24.69
Phase2 72.74 22.61 41.19 25.37
CP-BCS(Baseline) 17.78 21.50 16.89 11.92
After completing the training step, we use MalP as the test set (it is the only pseudocode summary dataset built for malware that we know until now). We compare Malsight to existing pseudocode summarization methods and the popular prompt-based general-purpose large model on dataset MalP. In the actual experiment, we built three versions of data for MalP: not-stripped, demi-stripped (only stripping the function name), and all-stripped (stripping the function name and all identifiers inside the function).
The closest version of the dataset to the real world, that is, the results of the all-stripped version as test set, are shown in Table III.
TABLE III: Comparison with baseline work
@@ 6.2.1 Code Summary Performance
As shown in Table III, we conducted experiments using evaluation methods such as BLEURT-sum, BLEU, ROUGE, and METEOR (the usability of BLEURT-sum is substantiated in the referenced paper) to carry out experiments on Malsight, WizardLM [57], and Code Llama [58], etc. We also measured model performance using criteria such as summary length, variance, and processing time, as these factors can directly impact algorithm-based evaluations (see Appendix I for details).
As previously mentioned, general-purpose large models, trained on extensive corpora, possess broader knowledge, providing an advantage in code summarization tasks. Among these, GPT, as a commercial large model, performed the best with a BLEURT-sum score of 0.6009. Malsight’s innovative fine-tuning approach offers a significant advantage in summarizing malware code, with its annotation generation effectively bridging the knowledge gap between specialized and general-purpose models. Malsight achieved the highest scores among all methods in both BLEURT-sum and METEOR, which we consider two of the most reliable evaluation metrics.
This conclusion can be corroborated by Figure 6. The figure shows that Malsight has better score evaluation results and less variance, i.e., more stable code summary output, on test results stripped of data.
Figure 6: Data distribution. Malsight presents a relatively stable high segmentation distribution, which beats GPT-3.5 work.
6.2.2 Annotation Extraction Effect
The annotation extraction model was trained and tested on the AnnoS dataset, and the test results are shown in the table The annotated extraction model was trained and tested on the AnnoS dataset to complete the sequence SL task, dividing the data into normal codes (represented by N-label), important function APIs (represented by A-label), and important strings (represented by S-label). The model test produces the confusion matrix shown in Table IV below.
TABLE IV: Annotation extraction confusion matrix
N-label(Actual) 1,847,647 9,345 11,721
A-label(Actual) 15,865 116,099 1,784
I-label(Actual) 21,639 1,688 38,933
The results show that the accuracy of the model is 96.99%, which is almost comparable to the results obtained by manual annotation.
6.3 RQ2. Ablation Experiment
To verify the usability of each module in Malsight, we conducted ablation experiments on the annotation module and the two phases of code summarization. We continued to use MalP as a test dataset, experimenting with different module combinations.
6.3.1 Module Ablation
As shown in Table V, the modules in Malsight complemented each other effectively. This ablation study was conducted on the not-stripped version of MalP to minimize result fluctuations due to the absence of dynamic annotation.
TABLE V: Ablation experiment
Model BLEURT-sum BLEU ROUGE-L METEOR
Full Model 66.29 10.95 26.97 18.80
w/o Annotation 60.05 7.65 22.76 14.93
w/o Phase 1 62.05 10.18 24.16 17.87
w/o Phase 1 and Annotation 56.48 7.35 21.99 13.99
w/o Phase 2 61.93 9.19 26.25 15.37
w/o Phase 2 and Annotation 59.13 7.42 26.72 13.32
TABLE VI: Performance on different levels of stripping
@@ 6.3.1 Module Ablation
We designed five experiments to test different combinations for the code summarization task, as follows: (1) Removing the annotation module and summarizing without any annotation, labeled as “w/o Annotation”; (2) Canceling the phase 1, labeled as “w/o Phase ”; (3) Canceling the phase 1 and removing the annotation module, labeled as “w/o Phase 1 and Annotatio”; (4) Canceling the phase 2, labeled as “w/o Phase 2”; (5) Canceling the phase 2 and removing the annotation module, labeled as “w/o Phase 2 and Annotation”.
@@
6.3.2 Annotater vs. Stripping
The influence of different levels of strip on the annotation module continues to be explored, and the results are shown in Table VI. Experiments have shown that a higher degree of stripping creates a tolerable performance degradation in the absence of annotation, demonstrating the robustness of Malsight. After the annotation module was ablated, it produced a 9.41% performance degradation (measured by a BLEURT-sum score) in the not-stripped test and nearly 12% in the all-stripped test, which proved that the annotator effectively resisted the adverse conditions caused by stripping.
Figure 7: Annotation Ablation. The Annotation generation module adds more semantic information to sentence A, making it able to output function functionality smoothly, but after ablation, it cannot.
In the generated results, we selected relatively representative sentences, as shown in Figure 7. In the yellow box, the output from the generation module without annotations shows that the large model is highly susceptible to biases due to hallucination [59]. Based on our observations, hallucinations can cause the code summary model to make false guesses about the types of parameters and internally called functions, leading to potential user misdirection. To mitigate this issue, we incorporate annotations, which not only enhance the model’s ability to summarize behavior, structure, and application scenarios but also reduce the nonsensical outputs caused by these illusions.
6.4 RQ3. Evaluation Algorithm Test
@@
Figure 8: Relevant method test results. The evaluation effect of two common algorithms and ML-based evaluation methods is not ideal, and their positive samples and negative samples show a certain degree of crossover.
The EvaS dataset was constructed to evaluate our code summary evaluation method alongside other popular methods. As previously mentioned, we use both positive and negative samples to test the effectiveness of these evaluation methods by assessing their ability to distinguish between the two. The performance of existing methods, illustrated in Figure 8, shows varying degrees of crossover between positive and negative samples for BLEU, METEOR, word2vec, and MoverScore. These four methods represent the current approaches in words’ overlap measure(BLEU and METEOR) and words’ embedding measure(word2vec and MoverScore), respectively.
@@
In Figure 8, we have established two types of decision boundaries: one being orthogonal to the X-axis (i.e., distinguishing between positive and negative samples by setting a threshold), and the other being a linear function forming a slanted line. The same process is represented by BLEURT-sum as shown in Figure 9, which shows the superiority of our method.
@@
Figure 9: BLEURT-sum performance. Compared with other methods, BLEURT-sum has significantly better ability to distinguish between positive and negative samples.
@@
Specifically, we calculated an F1-score for these six methods (ROUGE-L is not shown in the Figure 8) on the positive and negative sample classification task of code summary sentences to measure their ability. Our method achieved an F1-score exceeding 0.9999, significantly outperforming all other evaluation methods. Among the existing methods, METEOR performed the best with an F1-score of 0.9811, while BLEU had the lowest performance at only 0.85 as the lowest. None of the other methods achieved an F1-score above 0.95. Notably, our dataset was not meticulously curated; the negative samples were entirely random sentences with different meanings, which differ from real-world scenarios where the differences in meaning might be subtler. An F1-score below 95% may therefore indicate an unacceptable tendency to misclassify. This is because a 95% F1-score means that approximately 1 in 10 samples are incorrectly judged while upholding a high recall rate. As an evaluation method, the misjudgment can be further amplified by the model results evaluated using the evaluation scheme in the following works.
6.5 RQ4. Real World Experiments
6.5.1 Human Evaluation Experiments
So far, we have evaluated Malsight’s performance at the function level. However, this approach did not allow us to fully showcase the capabilities of the dynamic annotation module. Additionally, real-world malware predominantly manifests in executable files. Hence, we manually analyzed 10 real-world malware samples and selected 79 critical functions out of them, three experienced reverse engineers add summaries to these functions based on discussion. The complete workflow of Malsight was applied to these functions and the outputs was evaluated using both BLEURT-sum and human evaluation metrics. We further calculated the variance deviation of all calculated results as a reference to demonstrate the stability of our solution and the time required for code summarization for a single malware sample.
For the human evaluation metrics, we invited ten evaluators, including five reverse engineers who are rich in experience, three experienced, and two beginners. They were asked to focus on evaluating the usability of the code summarization results and provide a score of 0, 0.5 or 1, indicating whether the summary corresponded to the original code. It is noteworthy that due to the subjectivity of human evaluation of the summaries, different evaluators may have varying opinions on the summary of the same code. Therefore, we opted to evaluate usability only and did not solicit more detailed scores beyond 0,0.5 and 1, respectively corrrepond to usable, partially usable and unusable.
6.5.2 Performance From Different Evaluators

### v3

@@ V-A Evaluation Dataset Construction
Utilizing the code summary from the MalS dataset as a foundation for our research, we curate a positive and negative sample pair for the tuning of our new evaluation model BLEURT-sum. The positive sample consisted of two code summary result sentences sharing the same meaning is initialized as $\{S_{g},S_{r},1\}$, while the negative sample comprised two randomly different sentences and is initialized as $\{S_{g},S_{r},0\}$. $S_{g}$ and $S_{r}$ represent generated statements and reference statements, respectively.
@@
To build the dataset in $\{S_{g},S_{r},Score\}({Score}\in[0,1])$ format, one approach is to build an algorithm, $Score=GenSim(S_{g},S_{r},0or1)$, which automatically generate the $Score$ labels needed for model training. The key task is constructing a reliable $GenSim()$ function. When $S_{g}==SUM(f_{n})$, the $Score$ can be expressed using equation (14). This accounts for the fact that the SUM() function may produce different outputs for the same input, as it is a non-deterministic function.
@@
When constructing the $\{S_{g},S_{r}\}$ pairs, sentence structure dependencies are not considered, meaning that the score calculated in equation (14) is based purely on semantic information. To enhance this, we propose incorporating both sentence structure and semantics into the evaluation process. Our approach seeks to determine the appropriate balance between these two aspects in sentence similarity evaluation.
Taking the original 0,1 tag as the semantic feature, we further extract the static feature to get $s_{f}$. Assuming that the proportion of semantic information is $p$, the structural feature calculation function is $Struc()$, the following equations formalize this Score construction method.
$s_{f}=Struc(S_{g},S_{r})$ (15)
@@
V-B Evaluation Model
Based on our insights of the evaluation model, it should be able to combine both structural and semantic features and give an evaluation score for the specific task of code summary.
@@ V-B Evaluation Model
The rich training signals ensure the universality of BLEURT and the comprehensiveness of the evaluation angle. We then further fine-tune the BLEURT model on EvaS dataset. First, the generated text and the reference text are input together into the model and the vector is generated as shown in equation (17).
@@
In summary, existing word overlap methods often inflate scores due to common keywords like ”return” or ”initialize”, which may be incremental. Meanwhile, the BLEURT-sum improves performance by considering both word overlap and additional dimensions.
VI Experiments
TABLE II: Comparisons with baseline work
@@ VI Experiments
Our evaluation experiments are designed to answer the following four questions:
  1. 1.
    RQ1 (§ VI-B). How does Malsight performance compare across two training phases and mainstream code summarization models?
  2. 2.
    RQ2 (§ VI-C). How do module combinations and different stripping scenarios affect Malsight’s performance, particularly the annotator module?
  3. 3.
    RQ3 (§ VI-D). How does the new BLEURT-sum evaluation method compare to existing code summarization evaluation metrics?
  4. 4.
    RQ4 (§ VI-E). How does Malsight perform when applied to real-world malicious software, and how do assessments from human reverse engineers validate its usability?
VI-A Experimental Setup
Our experiments run on the Ubuntu 20.04 system, equipped with one Intel Xeon Silver 4210 CPU 2.20 GHz, and two NVIDIA A40 GPUs with 125 GB RAM. Binary file processing tools include Radare2, IDA Pro v7.5, GCC v9.4.0, and GNU Make v4.2.1. Our programming language is Python v3.8.13, with transformers v4.16.2, torch v2.1.2+cu121.
VI-B RQ1. Performance Test
VI-B1 Code Summary Performance
During training, the model undergoes two phases. The first phase involves fine-tuning malware source code (MalS) to learn malicious behavioral features. The second phase focuses on generating high-quality code summaries (BenignC) by incorporating semantic information from pseudocode and annotations. In both phases, we split the dataset into training, cross-validation, and test sets with a ratio of 7:2:1. We then evaluate the model at each stage to ensure its effectiveness. The result is shown in Table III.
TABLE III: Score During Two Train Phases
BLEURT-sum BLEU ROUGE-L METEOR
Phase1 74.17 18.06 38.56 24.69
Phase2 72.74 22.61 41.19 25.37
CP-BCS(Baseline) 17.78 21.50 16.89 11.92
TABLE IV: Performance comparisons on various levels of stripping
@@ VI-B1 Code Summary Performance
Since each phase uses a different dataset, their effectiveness should be verified independently rather than by direct score comparison. Considering there is no universally accepted threshold for model evaluation scores, we provide the performance of another state-of-the-art binary code summarization model (CP-BCS [55]) as a baseline. CP-BCS is tested on its own dataset, which consists of binaries compiled with GCC 7.3.0 for x86 architecture (32-bit) and then stripped. Our results in both phases demonstrate higher usability compared to one of the existing method CP-BCS.
Also, during the training process of transfer learning, we try different training data proportions and select the optimal proportion, as shown in Table V.
TABLE V: Performance in different training data proportion
Proportion BLEURT-sum BLEU ROUGE-L METEOR
1:1 59.12 8.98 23.04 16.14
1:2 59.81 9.15 23.50 16.42
3:4 59.53 8.68 22.95 16.39
4:3 62.14 9.80 25.11 16.88
* Limited by MalS, proportion 2:1 is unavailable.
* AVG summary length is measured in words.
After completing the training step, we use MalP as the test set (it is the only pseudocode summary dataset built for malware that we know until now). We compare Malsight to existing pseudocode summarization methods and prompt-based general-purpose large model on dataset MalP. Simultaneously, we built three versions of data for MalP: not-stripped, demi-stripped (only stripping the function name), and all-stripped (stripping all identifiers inside the function).
The closest version of the dataset to the real world, which is the results of the all-stripped version are shown in Table II. We conduct experiments using BLEURT-sum, BLEU, ROUGE, and METEOR (the usability of BLEURT-sum is substantiated in the referenced paper) to carry out experiments on Malsight, WizardLM [56], and Code Llama [57], etc. We also measure model performance using criteria such as summary length, variance, and processing time, as these factors can directly impact algorithm-based evaluations.
(a) Malsight vs Code Llama-7b
(b) Malsight vs WizardCoder-15b
Figure 6: Data distribution. The segmentation distribution yielded by the proposed Malsight is more stable than the ones of the previous code-Llama.
General-purpose large models, trained on extensive code corpora, offer an advantage in code summarization tasks, with Code-Llama performing the best, achieving a BLEURT-sum score of 56.00. Malsight’s innovative fine-tuning approach offers a significant advantage in summarizing malware code, with its annotation generation effectively bridging the knowledge gap between specialized and general-purpose models. It has been revealed that Malsight achieved the highest scores among all evaluation indexes.
This conclusion can be corroborated by Fig. 6. The figure shows that Malsight has better score evaluation results and less variance, i.e., more stable code summary output, on test results stripped of data.
VI-B2 Annotation Extraction Effect
The annotation extraction model is trained and tested on the AnnoS dataset, shown in Table VI. The annotated extraction model is trained and tested on the AnnoS dataset to complete the sequence SL task, dividing the data into normal codes (represented by N-label), important function APIs (represented by A-label), and important strings (represented by S-label). The model test produces the confusion matrix shown in Table VI below.
TABLE VI: Annotation extraction confusion matrix
N-label(Actual) 1,847,647 9,345 11,721
A-label(Actual) 15,865 116,099 1,784
I-label(Actual) 21,639 1,688 38,933
The results show that the accuracy of the model is 96.99%, which is comparable to the results obtained by manual annotation. Inevitably, some functions have ambiguous function names, or strings that appear to have certain characteristics, leading to negligible incorrect labeling.
VI-C RQ2. Ablation Study
To verify the usability of each module in Malsight, we conduct ablation experiments on the annotation module and the two phases of code summarization. Further, We use MalP as a test dataset, experimenting with different module combinations.
TABLE VII: Ablation experiment
Model BLEURT-sum BLEU ROUGE-L METEOR
Full Model 66.29 10.95 26.97 18.80
w/o Annotation 60.05 7.65 22.76 14.93
w/o Phase 1 62.05 10.18 24.16 17.87
w/o Phase 1 and Annotation 56.48 7.35 21.99 13.99
w/o Phase 2 61.93 9.19 26.25 15.37
w/o Phase 2 and Annotation 59.13 7.42 26.72 13.32
VI-C1 Module Ablation
As shown in Table VII, the modules in Malsight complement each other effectively. This ablation study is conducted on the not-stripped version of MalP to minimize result fluctuations due to the absence of dynamic annotation.
We conduct experiments under five settings: (1) Removing the annotation module and summarizing without any annotation, labelled as “w/o Annotation”; (2) Canceling phase 1, labelled as “w/o Phase ”; (3) Canceling phase 1 and removing the annotation module, labelled as “w/o Phase 1 and Annotation”; (4) Canceling the phase 2, labelled as “w/o Phase 2”; (5) Canceling the phase 2 and removing the annotation module, labelled as “w/o Phase 2 and Annotation”.
@@ VI-C1 Module Ablation
VI-C2 Annotater vs. Stripping
The influence of different levels of the strip on the annotation module continues to be explored, and the results are shown in Table IV. Experiments have shown that a higher degree of stripping creates a tolerable performance degradation in the absence of annotation, demonstrating the robustness of Malsight.
After the annotation module is ablated, it produces a 9.41% performance degradation (measured by a BLEURT-sum score) in the not-stripped test and nearly 12% in the all-stripped test, which has proved that the annotator effectively resists the adverse conditions caused by stripping.
Figure 7: Annotation ablation. The annotation generation module enriches sentence A with semantic information, enabling it to convey function functionality, which is lost after ablation.
In the generated results, we select relatively representative sentences, as shown in Fig. 7. In the yellow box, the output from the generation module without annotations shows that the large model is highly susceptible to biases due to hallucination [58]. Based on our observations, hallucinations can cause the code summary model to make false guesses about the types of parameters and internally called functions, leading to potential user misdirection. To mitigate this issue, we incorporate annotations, which not only enhance the model’s ability to summarize behavior, structure, and application scenarios but also reduce the nonsensical outputs caused by these illusions.
VI-D RQ3. Evaluation Algorithm Test
As previously mentioned, we use positive and negative samples to test the effectiveness of these evaluation methods by assessing their ability to distinguish between the two. The performance of existing methods, illustrated in Fig. 8, shows varying degrees of crossover between positive and negative samples for BLEU, METEOR, Word2vec, and MoverScore. These four methods represent the current approaches in words’ overlap measure(BLEU and METEOR) and words’ embedding measure(word2vec and MoverScore), respectively.
@@
Figure 8: Test results of popular text similarity algorithms. The evaluation effect of two common algorithms and ML-based evaluation methods is not ideal, and their positive samples and negative samples show a certain degree of crossover.
@@
In Fig. 8(a), we have established two types of decision boundaries: one being orthogonal to the X-axis (i.e., distinguishing between positive and negative samples by setting a threshold), and the other being a linear function forming a slanted line. The same process is represented by BLEURT-sum as shown in Fig. 9, which shows the superiority of our method.
Specifically, we calculate the F1-score for these six methods (ROUGE-L is not shown in Fig. 8) on the positive and negative sample classification task of code summary sentences to measure their ability. Our method achieves an F1-score exceeding 0.9999, significantly outperforming all other evaluation methods. Among the existing methods, METEOR performs best with an F1-score of 0.9811, while BLEU gets the lowest performance at only 0.85. None of the other techniques achieve an F1-score above 0.95.
(a) BLEURT-sum
(b) ROC Curve (Scaled)
@@
Figure 9: BLEURT-sum performance. BLEURT-sum has a significantly better ability to distinguish between positive and negative samples(a). Simultaneously, the ROC curve demonstrates a commendable AUC.(b)
@@
Notably, our dataset is not meticulously curated, the negative samples are entirely random sentences with different meanings, which differ from real-world scenarios where the differences in meaning might be subtler. An F1-score below 95% indicates a tendency to misclassify, with approximately 1 in 10 samples judged incorrectly. As an evaluation method, the misjudgment can be further amplified by the model results evaluated using the evaluation scheme in the following works.
VI-E RQ4. Real World Experiments
VI-E1 Human Evaluation Experiments
So far, we have evaluated Malsight’s performance at the function level, but this hasn’t fully demonstrated the dynamic annotation module’s capabilities. Since real-world malware predominantly appears as executables, we manually analyzed ten malware samples and selected 79 critical functions. Three experienced reverse engineers collaboratively provided summaries for these functions.
Malsight’s complete workflow is applied to these samples, with outputs evaluated using both BLEURT-sum and human evaluation metrics. We also calculate the variance deviation to assess the stability of the solution and measure the time required for summarizing a single malware sample.
For human evaluation, we invite ten evaluators, including five expert reverse engineers, three experienced engineers, and two beginners. They rate the summaries based on usability with scores of 0, 0.5, or 1, corresponding to unusable, partially usable, and usable, respectively. Given the subjectivity of human evaluation, we focus solely on usability rather than providing more detailed scores.
VI-E2 Performance From Different Evaluators